Circular No. 8211

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 S3 (LINEAR) M. Bezpalko, Lincoln Laboratory, Massachusetts Institute of Technol-

ogy, reports the LINEAR discovery of a comet, their Sept. 27 images showing a tail in p.a. 260^{o} (discovery observation below). Following web-posting on the NEO Confirmation Page, J. Young reported that CCD images taken with the 0.6-m reflector at Table Mountain on Sept. 28.4 and 29.4 UT reveal a 5′′ coma and a straight, narrow tail 20′′ long in p.a. 65^{o} and 57^{o}, respectively.

2003 UT  α_{2000}  δ_{2000}  Mag.
Sept.27.37818  4^{h}11^{m}01.^{s}82  +28^{o}54′42′′.2  20.1

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-S85.

T = 2003 Sept.19.814 TT  ω = 164.^{o}242
Ω = 226.322 }2000.0
q = 8.16889 AU  i = 151.586

SUPERNOVAE 2003gq, 2003hf, 2003hp, 2003ht, 2003hw, 2003ig A. V. Filippenko and R. Chornock, University of California, Berkeley,

report that inspection of CCD spectra (range 320–920 nm), obtained on Sept. 28 UT with the Keck-I 10-m telescope, reveals that SN 2003hf (IAUC 8182) is indeed of type II, as suggested on IAUC 8189; very strong, broad Hα emission is present, with essentially no absorption, but other features have P-Cyg profiles. Based on the available spectra, which closely resemble those of SN 1979C (Branch et al. 1981, Ap.J. 244, 780), it is likely that analysis of the light curve will show that the object is specifically of type II-L. SN 2003hp (IAUC 8188, 8189) is a 1998bw-like type-Ic supernova (see Foley et al. 2003, PASP 115, 1220, and references therein), roughly 1 month past maximum brightness. SN 2003ht (IAUC 8192) is of type II, in the nebular phase. SN 2003hw (IAUC 8198) is of type Ia, ∼ 2 months past maximum brightness. SN 2003ig (IAUC 8207) is of type Ic, within a week past maximum brightness. SN 2003gq (IAUC 8168), whose early-time spectrum (IAUC 8170) resembled that of SN 1991T but with low ejection velocities, continues to exhibit a spectrum very similar to that of SN 2002cx (Li et al. 2003, PASP 115, 453), characterized by unusually narrow lines.

2003 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green